The radius of a quarter circle is 3 meters. What is the quarter circle's perimeter?


Use 3.14 for ​𝜋.


1
Expert's answer
2021-12-02T18:32:58-0500

perimeter of a quarter circle of radius 'r' is, r + r + (2πr)/4 = 2r + πr/2.



Perimeter "=[2×3]+[\frac{3.14×3}{2}]"


Perimeter"=6+4.71=10.71meters"